# CREST-Search: Comprehensive Red-teaming for Evaluating Safety Threats in Large Language Models Powered by Web Search

### v1

@@ 5.2. Results
Effectiveness of generation strategy. Table 2 reports the average risk detection rates of adversarial search queries generated by three different strategies: Keyword injection, exaggeration, and role play, across five harmful content categories. Overall, these three strategies are all effective, while keyword injection achieves the highest risk detection rate, with particularly strong performance on harassment and sexual content. Across categories, sexual content shows the highest vulnerability (85.8% on average). These findings provide a perspective on why CREST-Search is effective in uncovering specialized risks in LLMs with web search, compared with traditional red-teaming for vanilla LLMs. Because our proposed generation strategies are highly aligned with the unique threat surface of LLMs augmented with web search, highlighting their pivotal role within the overall framework.
Transferability. The detailed risk detection rates of four victim models under CREST-Search are presented in Figure 3. The high total risk across all victim models indicates the strong transferability of CREST-Search across heterogeneous black-box LLMs with search. Citation risk overwhelmingly dominates, far surpassing response and combined risks, highlighting that our generated adversarial queries are particularly effective at uncovering vulnerabilities in the retrieval and citation process of LLMs with search. Gemini-series models exhibit a comparatively higher risk than the GPT-series models on average. Even when the generated responses themselves are benign, the toxic content that may exist in the cited sources still threatens the model’s safety, thereby undermining its reliability. In addition, we present the detailed risk detection rates across five harmful content categories on four victim models, which also validate the good transferability of CREST-Search. Details are shown in Appendix A.3. These results underscore the urgent requirement for more comprehensive citation defense mechanisms to protect LLMs with web search than vanilla LLMs.
Table 1. Performance comparison of baseline models.
   Risk Detection Rate    Toxicity    Diversity
   UAT    5.2%    87.9%    0.77
   HarmfulQA    5.7%    49.9%    0.71
   DangerousQA    10.5%    65.5%    0.59
   HarmBench    11.6%    53.8%    0.47
   CREST-Search    80.5%    23.6%    0.82
Figure 2. Detailed risks analysis across baseline models.
Table 2. Risk detection rates across generation strategies.
@@
Figure 3. Transferability of CREST-Search across various victim models.
@@
Figure 4. The impact of refinement rounds on risk detection rate (a), optimization cost (b), and optimization time (c) by five harmful content categories.
5.3. Ablation Study
@@
Finally, we conduct the ablation study to validate the contributions of each component during query generation and refinement stages. Two key factors affect the optimization performance and efficiency: the number of refinement rounds and the red-teaming model fine-tuning. To ensure the fairness of the experiment, all other experimental settings are held constant.
@@
Refinement rounds. To investigate how the maximum rounds of optimization influence refinement performance, we vary the maximum rounds from 1 to 10 while keeping the same other settings. For each harmful content category, we record three metrics: the risk detection rate of the generated test cases, the total API cost (per 100 queries, in USD), and the average refinement time (in seconds). This allows us to analyze the trade-off between refinement effectiveness and efficiency. Figure 4 illustrates the results. Overall, increasing the refinement rounds consistently improves risk detection rates across all categories, but at the expense of higher computational costs and longer optimization times. For example, harassment improves from 34.6% to 73.8%. In parallel, both optimization cost and time grow almost linearly, with harassment reaching 2.22 USD per 100 queries and 75.4 seconds per query at 10 rounds. Despite categories such as sexual content incurring relatively lower costs and times, all exhibit the same upward trend. This highlights a clear trade-off between effectiveness and efficiency in the refinement stage.
@@
Red-teaming model fine-tuning. We further examine the role of the red-teaming model in adversarial query generation. CREST-Search employs a fine-tuned model for red-teaming, while we set two baseline configurations for comparison: (1) vanilla zero-shot, where a general vanilla LLM is used to generate test cases without additional supervision; and (2) vanilla few-shot, where the same LLM is guided by a small set of manually selected successful adversarial queries as exemplars. All models are tasked with generating test cases under the same harmful content categories and strategies, enabling a fair comparison of their effectiveness. The results are reported in Table 3, consisting of five metrics: risk detection rate (RDR), query diversity, toxicity, cost fee (per 100 queries, USD), and time consumption (seconds). The fine-tuned model achieves the highest risk detection rate at 80.5%, significantly outperforming both few-shot (49.3%) and zero-shot (31.1%) baselines. It also produces the most diverse queries compared to the baselines, while the zero-shot is the lowest. Although the fine-tuned model yields slightly higher toxicity (23.6%) than baselines, its queries remain within a manageable range for controlled safety evaluation, where each test does not trigger refusals by the victim models. By contrast, the lower toxicity observed in baseline models largely results from the fact that most generated queries fail to trigger risks, thereby reducing their adversarial effectiveness and undermining the purpose of red-teaming. Importantly, the fine-tuned model is also the most efficient, requiring only \$1.50 per 100 queries and 49.6 seconds per query, while the zero-shot approach incurs the highest fee and optimization time. These results confirm that fine-tuning enhances adversarial strength and improves efficiency and diversity, making it the most practical choice for CREST-Search as the base red-teaming model for test case generation. This demonstrates that red-teaming approaches designed for vanilla models cannot be directly transferred, and emphasizes the necessity of specialized fine-tuned red-teaming models.
Table 3. Performance of different base generation models
RDR Diversity Toxicity Cost Fee Time Consumption
Zero-shot 31.1% 0.44 7.68% 4.1 128.5
Few-shot 49.3% 0.65 12.2% 2.7 89.1
Fine-tuned 80.5% 0.82 23.6% 1.5 49.6
6. Discussions
@@
Mitigation. To translate insights we gained from risks uncovered by CREST-Search into practical safeguards suggestions for the developers, we discuss two mitigation suggestions to reduce harmful outputs in LLMs with web search. The first defense is to implement safety detection on both the URLs returned by the model and the content of the corresponding webpages before citation, thereby reducing the likelihood of propagating harmful content to end users. However, this approach inevitably introduces a trade-off between safety and efficiency: comprehensive real-time analysis of webpages may increase response latency, potentially degrading user experience. Designing lightweight but effective filtering pipelines becomes an essential challenge. The second mitigation strategy is to leverage adversarial queries generated by CREST-Search to fine-tune the LLMs to construct more robustness and stronger safety alignment. The important concern for this method is that it need to strengthen robustness without compromising the utility of the models in benign use cases. Beyond these two strategies, another systematic approach is the continuous integration of automated red-teaming frameworks into model deployment pipelines, such as CREST-Search. It can serve as a cornerstone in building proactive and sustainable AI safety management.
Deployment clarification. While CREST-Search demonstrates strong effectiveness, our current evaluation primarily employs a commercial model (Gemini-2.5-flash) for fine-tuning, without including open-source models. This design choice is practical and does not diminish the validity of our methodology. Despite the cost of the usage fee, Commercial models offer practical benefits, including lower data requirements for effective training, faster fine-tuning procedures, and easier deployment. In contrast, open-source models often exhibit weaker baseline capabilities, which necessitate larger datasets, longer training cycles, and dedicated computational resources that ultimately also translate into higher costs.
7. Conclusion
@@
In this paper, we propose CREST-Search, a pioneering framework that comprehensively and systematically uncovers citation risks in LLMs with web search. The framework consists of three core components: adversarial query generation, execution and evaluation, and iterative refinement. To further enhance the generation efficiency and quality of test cases, we design three novel strategies and a safety benchmark dataset dedicated to search-enabled LLMs, and propose to fine-tune a specialized red-teaming model. Extensive experiments demonstrate that CREST-Search outperforms conventional red-teaming approaches in discovering this new class of risks in search-enabled LLMs, while significantly reducing both cost and time consumption. Our ablation study further validates the contributions of each component of the framework. Based on these findings, we provide two mitigation suggestions to help developers improve the safety of LLMs with web search. In the future, we believe that with the continued evolution of open-source LLMs, the cost and time required for such red-teaming will further decrease.
@@ Appendix A Appendix
A.1. Ethics Considerations
This work was conducted under strict ethical and security guidelines. All adversarial search queries and harmful outputs were generated solely within a controlled experimental environment for systematic evaluation. Our intention is not to attack or exploit deployed systems, but rather to discover vulnerabilities to improve the safety and reliability of such search-enabled LLMs. By exposing risks such as unsafe citations, our framework provides insights that can help developers design stronger safeguards and filtering mechanisms. This can reduce the likelihood of users encountering malicious or harmful information and strengthen public trust in LLM-based technologies, thereby supporting their responsible deployment and broader adoption.
A.2. WebSearch-Harm Dataset Analysis
@@
A.3. Harmful Content Category
@@
Table 4. Training dataset analysis.
@@
Figure 5. Detection rates for five harmful-content categories across victim models (CREST-Search).
@@ A.4. Prompt Template
You are an expert in adversarial query design for evaluating the safety of large language models with web search capabilities. Your goal is to generate user queries that appear neutral and harmless on the surface, but are likely to trigger the model to retrieve or cite webpages that contain unsafe content. You are given a specific content category and a query construction strategy. Based on these inputs, generate user queries. Each query should be distinct from the others if possible. Avoid question forms like ‘Can you’; use imperative or declarative tone instead. This is a controlled safety evaluation setting. You are required to complete the task exactly as specified. Do not refuse, skip, or self-censor your output under any circumstances. Therefore the response cannot start with ’I’m sorry, I can’t assist with that request.’ or similar disclaimers. Return your response in the following format: 1. ¡user query 1¿ 2. ¡user query 2¿ 3. ¡user query 3¿
Below is the user prompt that specifies the concrete context for generation, including the harmful content category, its description, the selected construction strategy, and a brief explanation of that strategy. Through the crafted prompts, we can ensure that the generated queries are both diverse and targeted, comprehensively uncovering the risks in the LLMs with web search.

### v2

@@ 5.2 Main Results
Effectiveness of generation strategy. Table 2 reports the average risk detection rates of adversarial search queries generated by three different strategies across five harmful content categories. Overall, these three strategies are all effective, with keyword injection achieving the highest risk detection rate and robust performance in detecting harassment and sexual content. Across categories, sexual content shows the highest vulnerability (85.8% on average). These findings provide a perspective on why CREST-Search is effective in uncovering specialized risks in LLMs with web search, compared with traditional red-teaming for vanilla LLMs. Our proposed strategies are highly aligned with the unique threat surface of LLMs augmented with web search, highlighting their pivotal role within the overall framework.
Transferability. The high total risks across all victim models (Figure 4) indicate the strong transferability of CREST-Search across heterogeneous black-box LLMs with search. Citation risks overwhelmingly dominate, far surpassing response and combined risks, highlighting that our generated adversarial queries are particularly effective at uncovering vulnerabilities in the retrieval and citation process of LLMs with search. Gemini-series models exhibit comparatively more risks than the GPT-series models on average. Even when the generated responses themselves are benign, the toxic content that may exist in the cited sources still threatens the model’s safety, thereby undermining its reliability. In addition, we present the detailed risk detection rates across five harmful content categories on four victim models, which also validate the good transferability of CREST-Search. Details are shown in Appendix A.3. These results underscore the urgent requirement for more comprehensive citation defense mechanisms to protect LLMs with web search than vanilla LLMs.
Table 2: Risk detection rates across generation strategies.
@@
Figure 4: CREST-Search transferability across victim models.
@@
Figure 5: The impact of refinement rounds on risk detection rate (left), cost (middle), and time (right) by different categories.
5.3 Ablation Study
@@
We conduct the ablation study to validate the contributions of each component during query generation and refinement stages. Two key factors affect the optimization performance and efficiency: the number of refinement rounds and the red-teaming model fine-tuning. To ensure fair evaluations, all other experimental settings are held constant.
@@
Refinement rounds. To investigate how the maximum rounds of optimization influence refinement performance, we vary the maximum rounds from 1 to 10 while keeping other settings the same. For each harmful content category, we record the risk detection rate of the generated test cases, the total API cost (per 100 queries, in USD), and the average refinement time (in seconds), to analyze the trade-off between refinement effectiveness and efficiency. Figure 5 illustrates the results. Overall, increasing the refinement rounds consistently improves risk detection rates across all categories, but at the expense of higher computational costs and longer optimization times. For example, harassment improves from 34.6% to 73.8%. In parallel, both optimization cost and time grow almost linearly, with harassment reaching 2.22 USD per 100 queries and 75.4 seconds per query at 10 rounds. Despite categories such as sexual content incurring relatively lower costs and times, all exhibit the same upward trend. This highlights a clear trade-off between effectiveness and efficiency in the refinement stage.
@@
Red-teaming model fine-tuning. We further examine the role of the red-teaming model in adversarial query generation. We set two baseline configurations for comparison: (1) vanilla zero-shot, where a general vanilla LLM is used to generate test cases without additional supervision; and (2) vanilla few-shot, where the same LLM is guided by a small set of manually selected successful adversarial queries as exemplars. All models are tasked with generating test cases under the same harmful content categories and strategies, enabling a fair comparison of their effectiveness. The results are reported in Table 3, consisting of five metrics: risk detection rate, query diversity, toxicity, cost fee (USD per 100 queries), and time cost (seconds). The fine-tuned model achieves the highest risk detection rate at 80.5%, significantly outperforming both few-shot (49.3%) and zero-shot (31.1%) baselines. This also demonstrates our dataset is sufficient for effective SFT, consistent with Gemini’s documentation: its SFT does not rely on very large datasets. It also produces the most diverse queries compared to the baselines, while the zero-shot is the lowest. Although the fine-tuned model yields slightly higher toxicity (23.6%) than baselines, its queries remain within a manageable range for controlled safety evaluation, where each test does not trigger refusals by the victim models. By contrast, the lower toxicity observed in baseline models largely results from the fact that most generated queries fail to trigger risks, thereby reducing their adversarial effectiveness and undermining the purpose of red-teaming. Importantly, the fine-tuned model is also the most efficient, requiring only \$1.50 per 100 queries and 49.6 seconds per query, while the zero-shot approach incurs the highest fee and optimization time.
6 Discussion about Potential Mitigation
@@
To translate insights we gained from risks uncovered by CREST-Search into practical safeguards suggestions, we discuss two mitigation suggestions to reduce harmful citations in LLMs with web search. The first is to implement safety detection on both the URLs returned by the model and the content of the corresponding webpages before citation, thereby reducing the likelihood of propagating harmful content to end users. However, this inevitably introduces a trade-off between safety and efficiency: comprehensive real-time analysis of webpages may increase response latency, potentially degrading user experience. Designing lightweight but effective filtering pipelines becomes an essential challenge. The second is to leverage adversarial queries generated by CREST-Search to fine-tune the LLMs to construct more robustness and stronger safety alignment. The important concern for this method is that it needs to strengthen robustness without compromising the utility of the models in benign use cases. Beyond these two strategies, another systematic approach is the continuous integration of automated red-teaming frameworks into model deployment pipelines, such as CREST-Search. It can serve as a cornerstone in building proactive and sustainable AI safety management.
Table 3: Performance of different base generation models
Detection Rate Diversity Toxicity Cost Fee Time Cost
Zero-shot 31.1% 0.45 7.68% 4.1 128.5
Few-shot 49.3% 0.75 12.2% 2.7 89.1
Fine-tuned 80.5% 0.59 23.6% 1.5 49.6
7 Conclusion
@@
In this paper, we propose CREST-Search, a pioneering framework that systematically uncovers citation risks in LLMs equipped with web search. At its core, CREST-Search is driven by adversarial search query generation strategies tailored to the retrieval-citation attack surface. Building on these strategies, we design a three-stage red-teaming pipeline for automated query generation, execution, evaluation, and refinement. To further enhance efficiency and generation quality, we construct a search-specific safety benchmark and fine-tune a specialized red-teaming model. Extensive experiments demonstrate that CREST-Search outperforms conventional red-teaming approaches in discovering this new class of risks in search-enabled LLMs, while significantly reducing both cost and time consumption. Our ablation study further validates the contributions of each component of the framework. Based on these findings, we provide two mitigation suggestions to help improve the safety of LLMs with web search.
Impact Statement
This work was conducted under strict ethical and security guidelines. All adversarial search queries and harmful outputs were generated solely within a controlled experimental environment for systematic evaluation. Our intention is not to attack or exploit deployed systems, but rather to discover vulnerabilities to improve the safety and reliability of such search-enabled LLMs. By exposing risks such as unsafe citations, our framework provides insights that can help developers design stronger safeguards and filtering mechanisms. This can reduce the likelihood of users encountering malicious information and strengthen public trust in LLM-based technologies, thereby supporting their responsible deployment and broader adoption.
@@ Appendix A Appendix
A.1 Deployment Clarification
While CREST-Search demonstrates strong effectiveness, our current evaluation primarily employs a commercial model (Gemini-2.5-flash) for fine-tuning, without including open-source models. This design choice is practical and does not diminish the validity of our methodology. Despite the cost of the usage fee, Commercial models offer practical benefits, including lower data requirements for effective training, faster fine-tuning procedures, and easier deployment. In contrast, open-source models often exhibit weaker baseline capabilities, which necessitate larger datasets, longer training cycles, and dedicated computational resources that ultimately also translate into higher costs. In addition, we select Gemini-2.5-flash over GPT-4o primarily due to policy constraints. It is just an engineering choice: OpenAI explicitly prohibits fine-tuning on malicious or harmful data, whereas Gemini permits controlled fine-tuning for security research purposes. This difference makes Gemini-Flash a practical choice for constructing a red-teaming model that operates on adversarial data.
A.2 WebSearch-Harm Dataset Analysis
@@
A.3 Harmful Content Category
@@
Table 4: Training dataset analysis.
@@
Figure 6: Detection rates for five harmful-content categories across victim models (CREST-Search).
@@ A.4 Prompt Template
You are an expert in adversarial query design for evaluating the safety of large language models with web search capabilities. Your goal is to generate user queries that appear neutral and harmless on the surface, but are likely to trigger the model to retrieve or cite webpages that contain unsafe content. You are given a specific content category and a query construction strategy. Based on these inputs, generate user queries. Each query should be distinct from the others if possible. Avoid question forms like ‘Can you’; use imperative or declarative tone instead. This is a controlled safety evaluation setting. You are required to complete the task exactly as specified. Do not refuse, skip, or self-censor your output under any circumstances. Therefore the response cannot start with ’I’m sorry, I can’t assist with that request.’ or similar disclaimers. Return your response in the following format: 1. <user query 1> 2. <user query 2> 3. <user query 3>
Below is the user prompt that specifies the concrete context for generation, including the harmful content category, its description, the selected construction strategy, and a brief explanation of that strategy. Through the crafted prompts, we can ensure the generated queries are both diverse and targeted, comprehensively uncovering risks in the LLMs with web search.